Sol Summary #5 – February 15th

Sol: 5

Summary Title: Geology, Biology, and EVA-ology

Author’s name: Nathan Hadland

Mission Status: Nominal

Sol Activity Summary: The weather on Mars this morning was much more pleasant than on Sol 4. The sky was clear, but the muddy conditions persisted until late morning when the heat of the day started to clear the landscape of its saturation. There was no EVA today due to the unsafe ground conditions, so the sol was spent catching up on paperwork, characterizing regolith samples in the Science Dome, cleaning, and working in the GreenHab. The ISRU project reached a pivotal point today, with the Arabidopsis thaliana germinations completed and the seedlings transferred to regolith samples collected on previous EVAs using an agar media plug. Those experiments are now in the GreenHab and will stay there for the duration of the experiment. The team also made headway on our EVA optimization project. We finished designing our quantitative study to evaluate the effectiveness of EVAs and will be applying the measure to upcoming EVAs. We also worked on designing pre and post check lists for future crews to make EVAs safer, more efficient, and easier which will be available at the end of the mission. The crew’s journalist worked all day on her media and outreach project, interviewing each crewmember about their role, and recording while everyone was working. She is also designing and outreach video for kids. After the communications window and an Irish meal of shepherd’s pie tonight, the team will have our nightly meeting and leadership handover to discuss tomorrow.

Look Ahead Plan: Weather permitting tomorrow, the crew will go on EVA to continue geological sampling and testing of EVA checklists, procedures, and quantitative measures of effectiveness. Otherwise, we will continue refining the SOPs and other documentation to improve the workflow of future crews and continue to work in the laboratory. Tomorrow will also see another rotation of Commander and Executive Officer, so a new day will bring new leadership.

Anomalies in work: None

Weather: Partly cloudy with high winds throughout the day.

Crew Physical Status: Nominal

EVA: None

Reports to be filed: Sol Summary, EVA request, Operations Report, GreenHab Report, Journalist Report, Astronomy Report, HSO Report

Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site